OpenStudy (anonymous):

solve the equation log base 4 (x^2-3x)=1

OpenStudy (anonymous):

\[1 = \log_{4} 4\]

OpenStudy (anonymous):

but how did you get that answer?

OpenStudy (anonymous):

you can transform the right side into logarithms\[\log_{4} \left( x^2 - 3x \right) = \log_{4} 4\]and then you get\[\left( x^2 - 3x \right) = 4\]

OpenStudy (anonymous):

x^2-3x = 4 Can you continue?